OpenStudy (anonymous):

what do you do when you have sin2x+sinx=0?

OpenStudy (anonymous):

Are they asking if this is an identity?

OpenStudy (anonymous):

No, it says solve the equation for solutions [0, 2pie).

OpenStudy (anonymous):

They must want you to plug in what you have for x into that equation, because this is not an identity.

OpenStudy (anonymous):

No, the x is just the variable they use, you really don't need it in the equation. It can be read as sin2+sin=0.

OpenStudy (anonymous):

hmm I just finished doing something like this in calculus.

OpenStudy (anonymous):

Lol, im in trig. And this part is on my final.

OpenStudy (anonymous):

I see.. I just took my final! DO you mind posting a snip of this question?

OpenStudy (anonymous):

sin2x+sinx=0 2sinxcosx+sinx=0 sin x(2cosx+1)=0 either sinx=0=sin2npi, x=2npi,where n=........,-3,-2,-1,0,1,2,3,........ or 2cosx+1=0,2cosx=-1,cosx=-1/2=-cos60=cos(180-60),cos(180+60) x=120,240

OpenStudy (anonymous):



OpenStudy (anonymous):

we get infinite solutions of x by adding 360k from 120 and240,where k=.....,-2,-1,0,1,2,..

OpenStudy (anonymous):

What about \[\cos2x=\sqrt{2} - \cos2x\]

OpenStudy (anonymous):

sin2x+sinx=0 means we have to solve for x for infinite solutions which i have done

OpenStudy (anonymous):

\[\cos 2x=\sqrt{2}-\cos 2x\] \[\cos 2x+\cos2x=\sqrt{2}\] \[2\cos2x=\sqrt{2}\] \[\cos 2x=\frac{ \sqrt{2} }{ 2 }=\frac{ 1 }{\sqrt{2} }=\cos \frac{ \pi }{ 4 },\cos \frac{ -\pi }{4 }\] \[2x=\frac{ \pi }{4 },\frac{ -\pi }{ 4 }=\frac{ \pi }{ 4 }\pm2\pi ,\frac{ -\pi }{ 4 }\pm2\pi \] \[x=\frac{ \pi }{ 8 }\pm \pi,\frac{ -\pi }{ 8 }\pm \pi \]